# Xiaofeng Yang, Yuhui Zhang, Yifeng Zheng, Xiuqi Huang, Minfeng Zhuparallel-$\parallel$, Yuxin Ma, Bin Cui§, Wei Chen

### v2

@@ IV-B Knowledge Organization
As depicted in Figure 5, the knowledge graph adopts a tree-based structure for knowledge organization. The nodes $\{\mathcal{V}\}$ are structured into five primary types: database, table, column, value, and jargon, each comprising various components (e.g., description, usage) and uniquely identified by a name. To address the common challenge of terminological inconsistencies in user queries (e.g., synonyms, acronyms), an additional node type, alias, has been introduced. This node type contains alternative terms associated with the official name of other node types and can be dynamically updated in real-world applications. The relationships between these nodes are represented by edges ${}\{\mathcal{E}\}$, which delineate both logical relationships among the primary node types and associative relationships between alias nodes and other primary nodes.
To facilitate efficient knowledge retrieval, we develop a task-aware indexing mechanism for graph nodes, utilizing Elasticsearch [39] for full-text search and StarRocks [40] for embedding search. This supports both lexical and semantic matching of knowledge nodes in response to user queries. The indexing structure is designed as triplets ($\{name,content,tag\}$), where the $content$ field is a concatenation of knowledge components specified based on the various requirements of downstream tasks. For instance, some tasks necessitate the calculation logics while others only need descriptions for successful completion. By dynamically selecting the appropriate index, we ensure that knowledge retrieval is both efficient and effective.
@@ IV-C Knowledge Utilization
3:  $\mathcal{V_{Q}}\leftarrow lex\_search(\mathcal{Q},\mathcal{G})+sem\_search(\mathcal{Q},\mathcal{G})$
@@
6:     if $v_{i}.type==alias$ then
7:        $v_{i}\leftarrow backtrack(v_{i})$ // Backtrack to a primary node
@@
10:     $score_{i}\leftarrow\omega_{1}\cdot lex\_eval(\mathcal{Q},v_{i})+\omega_{2}\cdot sem\_eval(\mathcal{Q},v_{i})+$          $\omega_{3}\cdot LLM\_eval(\mathcal{Q},v_{i})$
@@
12:  $\mathcal{V_{Q}}.sort(score_{i})$ // Rank by matching score
13:  return  $\mathcal{V_{Q}}.topK$
@@
DSL Translation. The final step involves translating the query into a DSL specification, a common routine in BI scenarios [10]. This JSON structure specifies the relevant data and processing requirements, including fields such as ‘MeasureList’ (i.e., numerical columns), ‘DimensionList’ (i.e., categorical columns), and ‘ConditionList’ (i.e., filters). We prompt an LLM for DSL translation, providing detailed instructions and in-context examples to improve its performance. The generated DSL specification is validated using JSON Schema [42] to ensure syntactic and semantic correctness. This specification can then be directly converted to high-level languages like SQL and Vega-Lite based on predefined rules, or be used to enhance free-form code generation for complex tasks like NL2Insight, thereby facilitating LLMs’ performance in practical business settings.
@@ V Inter-Agent Communication
In this section, we introduce DataLab’s Inter-Agent Communication module, which facilitates efficient communication among multiple agents to complete complex BI tasks.
@@
Information Format Structure. A critical consideration in multi-agent collaboration is what ‘language’ agents use to communicate. In BI scenarios, the information exchanged among agents is diverse, encompassing types such as SQL queries, Python code, and charts. This variety makes traditional multi-agent frameworks [16, 17] that rely on natural language for information sharing inadequate. To address this, we design a structured information format tailored to the unique characteristics of BI scenarios.
@@
Figure 6: Workflow of Inter-Agent Communication.
Information Sharing Protocol. Another important question to consider is how to ensure efficient information sharing among agents. One extreme approach is that each agent receives information solely from its predecessor based on a plan chain [43]. While this minimizes data volume, it may cause the agent overlooking essential context (e.g., background information pertinent to the task). Conversely, allowing any agent unrestricted access to the information of all others is inefficient, as irrelevant context can degrade LLMs’ reasoning quality [44] and introduces additional computational overhead [45] during inference. Therefore, each agent should only receive the most relevant information for task completion.
@@ VI Cell-based Context Management
Thanks to the Inter-Agent Communication module (Section V), DataLab can handle complex BI tasks that require the collaboration of multiple agents. Specifically, user queries will trigger the presentation or modification of cells in the notebook, each corresponding to specific information units in the shared buffer. However, the previous module primarily aims at facilitating individual task completion for single data roles. In contrast, real-world BI scenarios often involve multiple data roles working on different tasks and collaborating within a unified notebook. This typically results in a multitude of multi-language cells (e.g., SQL, Python, Chart) generated or altered by either agents or users themselves. When addressing a new user query, it is crucial to efficiently provide the proxy agent with the necessary contexts from the notebook. Simply supplying all cells and their associated information units is impractical due to inefficiency and high token costs. Therefore, we aim to identify the minimum set of relevant cells to minimize token usage without compromising agent performance. To achieve this, we model cell dependencies within the notebook as a DAG based on variable references, and propose an adaptive context retrieval mechanism.
@@
1:  $v\_hash,cell\_refs\leftarrow\emptyset,\emptyset$
@@
4:     if $c.type==Python$ then
5:        $ast\leftarrow construct\_ast(c)$
6:        $new\_v\leftarrow find\_global\_variables(ast)$
7:        $v\_hash[new\_v]\leftarrow c$
8:     else if $c.type==SQL$ then
9:        $data\_v\leftarrow find\_data\_variable(c)$
10:        $v\_hash[data\_v]\leftarrow c$
@@
15:     $external\_v\leftarrow find\_external\_variables(c)$
16:     $cell\_refs[c]\leftarrow find\_ref\_cells(external\_v,v\_hash)$
@@
18:  return  $\mathcal{G}\leftarrow construct\_dag(cell\_refs)$
@@
  • Identify new variables. For Python cells, we construct an abstract syntax tree (AST) from the raw code to find global variables that are accessible across the entire notebook (e.g., function/class definitions, package imports). We exclude local variables as they are only visible within their scope. For SQL cells, any SELECT’s output is stored in a data variable (e.g., DataFrame) for future use, and thus represents a new variable. Conversely, Markdown and Chart cells do not produce variables that can be referenced elsewhere, and are therefore omitted. We store the variable-cell associations using a hash table.
  • Find referenced cells. Based on the hash table, we locate each cell’s referenced cells by identifying its external variables that are defined in other cells. For Python and SQL cells, this can be easily achieved with ASTs. For Chart cells, the underlying data variable serves as the reference point. As Markdown cells do not associate with any variables, they are excluded from this step. Consequently, a DAG of the notebook can be constructed using the extracted cell references.
@@ VII Experiment
In this section, we evaluate DataLab’s performance using both public research benchmarks and proprietary business datasets from Tencent, including End-to-End Performance Comparison, Sensitivity Analysis, Domain Knowledge Incorporation Evaluation, Inter-Agent Communication Evaluation, and Cell-based Context Management Evaluation.
@@
Data Preparation NL2SQL Spider [47] Execution Accuracy DataLab (Ours) DAIL-SQL [48] DIN-SQL [49]
80.70 -2.54% 83.60 82.80
BIRD [50] Execution Accuracy DataLab (Ours) DAIL-SQL DIN-SQL
61.33 +9.71% 57.41 55.90
NL2DSCode DS-1000 [4] Pass Rate DataLab (Ours) CoML [51] Code Interpreter [52]
53.80 +21.72% 44.20 51.60
DSEval [53] Pass Rate DataLab (Ours) CoML Code Interpreter
80.99 +12.64% 71.90 80.58
Data Analysis NL2Insight DABench [54] Accuracy DataLab (Ours) AutoGen [15] AgentPoirot [32]
75.10 +5.06% 71.48 75.88
InsightBench [32] LLaMA-3-Eval DataLab (Ours) AutoGen AgentPoirot
0.37 +19.35% 0.31 0.35
ROUGE-1 DataLab (Ours) AutoGen AgentPoirot
0.33 +17.86% 0.28 0.35
Data Visualization NL2VIS nvBench [55] Execution Accuracy DataLab (Ours) LIDA [56] Chat2Vis [57]
53.90 +0.13% 54.71 53.83
VisEval [58] Pass Rate DataLab (Ours) LIDA Chat2Vis
75.99 +5.67% 74.66 71.91
Readability Score DataLab (Ours) LIDA Chat2Vis
3.73 +0.81% 3.77 3.70
  • NOTE: Percentage gain is calculated relative to the weaker SOTA baseline marked with underlines.
@@ VII-A End-to-End Performance
To demonstrate the capabilities of DataLab as a unified BI platform, we first compare its end-to-end performance with SOTA LLM-based baselines on four typical BI tasks. We utilize GPT-4 [59] as the foundation model for all methods.
@@ VII-A1 Settings
The NL2SQL task converts natural language to SQL queries, typically marking the start of a BI workflow. We use two benchmarks (i.e., Spider [47] and BIRD [50]) and compare with two baselines (i.e., DAIL-SQL [48] and DIN-SQL [49]). We use the Execution Accuracy (EX) as the evaluation metric, which measures the execution equivalence of the generated SQL queries with the ground truth.
The NL2DSCode task converts natural language to data science code using Python libraries like NumPy and Pandas, which happens frequently throughout the BI workflow. We use two benchmarks (i.e., DS-1000 [4] and DSEval  [^1] We only evaluate DSEval-LeetCode and -SO due to implementation issues. [53]) and compare with two baselines (i.e., CoML [51] and Code Interpreter [52]). Pass Rate is used as the evaluation metric, which divides the number of passed problems by all problems.
The NL2VIS task converts natural language to data visualizations based on either Python libraries like Matplotlib or visualization grammars like Vega-Lite. We use two benchmarks (i.e., nvBench [55] and VisEval [58]) and compare with two baselines (i.e., LIDA [56] and Chat2Vis [57] [^2] As the two baselines lack support for NL queries related to multiple data tables, we only evaluate on single-table queries for fair comparison.. For nvBench, we use the EX metric for evaluation, which measures the equivalence of the generated visualizations with the ground truth based on the presented data values and chart types [60]. For VisEval, we use the Pass Rate metric to measure the ratio of valid or legal results divided by all queries, and the Readability Score judged by GPT-4V(ision) [61] to measure the overall quality of the generated visualizations [58].
@@ VII-A2 Results
As shown in Table I, DataLab achieves comparable performance or even surpass the SOTA LLM-based baselines (most of which only focus on a single task) on all four BI tasks. Specifically, DataLab outperforms all baselines on benchmarks including BIRD, DS-1000, DSEval, InsightBench, and VisEval, spreading over each critical BI stage. This demonstrates the superiority of DataLab in unifying the BI workflow with a single LLM-based agent framework.
For tasks that require the generation of symbolic languages (e.g., NL2SQL, NL2DSCode, NL2VIS), DataLab consistently performs well primarily due to the intermediate DSL specifications generated by our Domain Knowledge Incorporation module. Although most research benchmarks lack the necessary information for extracting table/column knowledge, DataLab adopts a meticulously-designed data profiling strategy as an alternative (Section IV-C) to fully utilize the provided data schema, enabling LLMs to correctly identify and associate the semantic relationships between data columns and NL queries, which are crucial to generate high-quality DSLs. Consequently, compared to merely feeding the original pure NL queries, these intermediate DSLs can significantly improve LLM-based agents’ performance on generating higher-level languages like SQL queries, Python code, or Vega-Lite specifications, as also shown in previous works [62].
@@ VII-C2 Downstream Tasks
TABLE II: Ablation study on Domain Knowledge Incorporation
Task / Metric S1 S2 S3
Schema Linking / Recall @5 (%) 41.02 71.79 79.49
NL2DSL / Accuracy (%) 32.52 61.66 91.10
@@ VII-E1 DAG Construction
VII-E2 Task Completion
As illustrated in Table IV, DataLab achieves a satisfactory Accuracy under both settings. Further analysis reveals that certain Markdown cells may contain critical information for task completion, which are occasionally failed to retrieve by our context retrieval mechanism due to limitations of embedding similarity [68]. This accounts for the slight 4.67% drop in Accuracy under S2. Conversely, S2 significantly reduces the Token Cost per Query by 61.65% compared to S1. This is achieved by identifying the minimum set of relevant cells based on DAGs. These results demonstrate the cost-effectiveness of this module while maintaining acceptable performance.
VIII Related Work
Business Intelligence Platforms. BI platforms support users in analyzing business data for decision-making [8, 12]. Representatives like Tableau [69], Power BI [70], and Databricks [71] provide GUIs to support user interactions for data transformation and dashboard generation. These platforms also integrate natural language interfaces [72, 73] to lower the burden of manual operation. Quamar et al. [1] proposed an ontology-based method based on business models to provide semantic information and reasoning capability for query interpretation. The emergence of LLMs has further enhanced the domain knowledge integration and visualization generation abilities of BI platforms [74, 75]. Compared with these existing tools that mostly target data analysis scenarios, we go beyond to support more comprehensive stages (e.g., data collection and preparation) and roles (e.g., data engineers and scientists) in BI scenarios, which significantly reduces the cost of platform switch and requirement communication.

### v3

@@ IV-B Knowledge Organization
As depicted in Figure 4, the knowledge graph adopts a tree-based structure for knowledge organization. The nodes $\{\mathcal{V}\}$ are structured into five primary types: database, table, column, value, and jargon, each comprising various components (e.g., description, usage) and uniquely identified by a name. To address the common challenge of terminological inconsistencies in user queries (e.g., synonyms, acronyms), an additional node type, alias, has been introduced. This node type contains alternative terms associated with the official name of other node types. It is predetermined based on enterprise-specific glossaries and can be dynamically updated in real-world applications. The relationships between these nodes are represented by edges ${}\{\mathcal{E}\}$, which delineate both logical relationships among the primary node types and associative relationships between alias nodes and other primary nodes.
[…]
@@ IV-C Knowledge Utilization
3:  $\mathcal{V_{Q}}\leftarrow\text{lex\_search}(\mathcal{Q},\mathcal{G})+\text{sem\_search}(\mathcal{Q},\mathcal{G})$
@@
6:     if $v_{i}.\text{type}==\text{alias}$ then
7:        $v_{i}\leftarrow\text{backtrack}(v_{i})$ // Backtrack to a primary node
@@
10:     $\text{score}_{i}\leftarrow\omega_{1}\cdot\text{lex\_eval}(\mathcal{Q},v_{i})+\omega_{2}\cdot\text{sem\_eval}(\mathcal{Q},v_{i})+$          $\omega_{3}\cdot\text{LLM\_eval}(\mathcal{Q},v_{i})$
@@
12:  $\mathcal{V_{Q}}.\text{sort}(\text{score}_{i})$ // Rank by matching score
13:  return  $\mathcal{V_{Q}}.\text{topK}$
@@
DSL Translation. The final step translates the query into a DSL specification, a common routine in BI scenarios [10]. This JSON structure specifies the relevant data and processing requirements, including fields such as ‘MeasureList’ (i.e., numerical columns), ‘DimensionList’ (i.e., categorical columns), and ‘ConditionList’ (i.e., filters). We prompt an LLM for DSL translation, providing detailed instructions and in-context examples to improve its performance. The generated DSL specification is validated using JSON Schema [37] to ensure syntactic and semantic correctness. This specification can then be directly converted to high-level languages like SQL and Vega-Lite based on predefined rules, or used to enhance free-form code generation for complex tasks like NL2Insight, thereby facilitating LLMs’ performance in business settings.
@@ V Inter-Agent Communication
In this section, we introduce DataLab’s Inter-Agent Communication module, which facilitates efficient communication among multiple agents to complete complex BI tasks. These agents are created and optimized beforehand through DAG-based workflows (Section III) to meet different user needs. For example, a SQL writing agent is specialized for NL2SQL tasks and a chart generation agent is for NL2VIS tasks.
@@
Information Format Structure. A critical consideration in multi-agent collaboration is what ‘language’ agents use to communicate. In BI scenarios, the information exchanged among agents is diverse, encompassing types such as SQL queries, Python code, and charts. This variety poses a significant challenge in ensuring accurate information sharing without introducing redundancy or miscommunication. Existing frameworks [17, 18] that rely on unstructured natural language for communication suffer from vagueness and inefficiencies. To address this, we design a structured information format [19] tailored to unique characteristics of BI scenarios.
Figure 5: Workflow of Inter-Agent Communication.
@@
Information Sharing Protocol. Another key challenge to resolve is how to ensure efficient information sharing among agents. One extreme approach is that each agent receives information solely from its predecessor based on a plan chain [38]. While this minimizes data volume, it may cause the agent overlooking essential context (e.g., background information pertinent to the task). Conversely, allowing any agent unrestricted access to the information of all others is inefficient, as irrelevant context can degrade LLMs’ reasoning quality [39] and introduces additional computational overhead [40] during inference. Therefore, each agent should only receive the most relevant information for task completion.
@@ VI Cell-based Context Management
Thanks to the Inter-Agent Communication module (Section V), DataLab can handle complex BI tasks that require the collaboration of multiple agents. Specifically, user queries will trigger the presentation or modification of cells in the notebook, each corresponding to specific information units in the shared buffer. However, the previous module primarily aims at facilitating individual task completion for single data roles. In contrast, real-world BI scenarios often involve multiple data roles working on different tasks and collaborating within a unified notebook. This typically results in a multitude of multi-language cells (e.g., SQL, Python, Chart) generated or altered by either agents or users themselves. When addressing a new user query, it is crucial to efficiently provide the proxy agent with the necessary contexts from the notebook. Simply supplying all cells and their associated information units is impractical due to inefficiency and high token costs. Therefore, we aim to identify the minimum set of relevant cells to minimize token usage without compromising agent performance. However, notebooks are inherently dynamic and collaborative environments, where iterative modifications reshape contextual relevance, and multi-role contributions create fragmented contexts requiring task-specific filtering. To address this challenge, we model cell dependencies within the notebook as a DAG based on variable references, and propose an adaptive context retrieval mechanism.
@@
[…]
@@
4:     if $c.\text{type}==\text{Python}$ then
[…]
8:     else if $c.\text{type}==\text{SQL}$ then
[…]
@@
[…]
@@
[…]
@@
  • Identify new variables. For Python cells, we construct an abstract syntax tree (AST) to find global variables accessible throughout the notebook (e.g., function/class definitions, package imports). We exclude local variables as they are only visible within their scope. For SQL cells, any SELECT’s output is stored in a data variable (e.g., a DataFrame) for future use, and thus represents a new variable. Markdown and Chart cells do not produce variables that can be referenced elsewhere, and are thus omitted. We store the variable-cell associations using a hash table.
  • Find referenced cells. Based on the hash table, we locate each cell’s referenced cells by identifying its external variables defined in other cells. For Python and SQL cells, this can be easily achieved with ASTs. For Chart cells, the underlying data variable serves as the reference point. As Markdown cells do not associate with any variables, they are excluded from this step. Using the extracted cell references, a DAG of the notebook can be constructed.
@@ VII Experiment
We evaluate DataLab on both public research benchmarks (§A, §B) and proprietary datasets from Tencent (§C, §D, §E).
@@
Data Preparation NL2SQL Spider [42] Execution Accuracy DataLab (Ours) DAIL-SQL [25] PURPLE [21] CHESS [29]
80.70 83.60 87.80 87.20
BIRD [43] Execution Accuracy DataLab (Ours) DAIL-SQL [25] PURPLE CHESS
61.33 57.41 68.12 68.31
NL2DSCode DS-1000 [4] Pass Rate DataLab (Ours) CoML [44] Code Interpreter [45] Open Interpreter [46]
53.80 44.20 51.60 50.50
DSEval [47] Pass Rate DataLab (Ours) CoML Code Interpreter Open Interpreter
80.99 71.90 80.58 78.10
Data Analysis NL2Insight DABench [48] Accuracy DataLab (Ours) AutoGen [16] AgentPoirot [26] -
75.10 71.48 75.88 -
InsightBench [26] LLaMA-3-Eval DataLab (Ours) AutoGen AgentPoirot -
0.37 0.31 0.35 -
ROUGE-1 DataLab (Ours) AutoGen AgentPoirot -
0.33 0.28 0.35 -
Data Visualization NL2VIS nvBench [49] Execution Accuracy DataLab (Ours) LIDA [30] Chat2Vis [27] CoML4VIS [50]
53.90 54.71 53.83 51.12
VisEval [50] Pass Rate DataLab (Ours) LIDA Chat2Vis CoML4VIS
75.99 74.66 71.91 75.27
Readability Score DataLab (Ours) LIDA Chat2Vis CoML4VIS
3.73 3.77 3.70 3.80
  • NOTE: The best method is marked in bold, while the second-best method is marked with underlines.
@@ VII-A End-to-End Performance
To demonstrate the capabilities of DataLab as a unified BI platform, we first compare its end-to-end performance with SOTA LLM-based baselines on four typical BI tasks. For fair comparison, the baselines employed are all prompt- or agent-based methods without any pre-training or supervised fine-tuning, similar to DataLab’s agent framework. We utilize GPT-4 [51] as the foundation model for all methods.
@@ VII-A1 Settings
The NL2SQL task converts natural language to SQL queries, typically marking the start of a BI workflow. We use two benchmarks (i.e., Spider [42] and BIRD [43]) and compare with three baselines (i.e., DAIL-SQL [25], PURPLE [21], and CHESS [29]). We use Execution Accuracy (EX) as the evaluation metric, which measures the execution equivalence of the generated SQL queries with ground truth.
The NL2DSCode task converts natural language to data science code using Python libraries like NumPy and Pandas, which happens frequently throughout the BI workflow. We use two benchmarks (i.e., DS-1000 [4] and DSEval  [^1] We only evaluate DSEval-LeetCode and -SO due to implementation issues. [47]) and compare with three baselines (i.e., CoML [44], Code Interpreter [45], and Open Interpreter [46]). Pass Rate is used as the evaluation metric, which divides the number of passed problems by all problems.
The NL2VIS task converts natural language to data visualizations based on either Python libraries like Matplotlib or visualization grammars like Vega-Lite. We use two benchmarks (i.e., nvBench [49] and VisEval [50]) and compare with three baselines (i.e., LIDA [30], Chat2Vis [27], and CoML4VIS [50] [^2] As some baselines lack support for NL queries related to multiple data tables, we only evaluate on single-table queries for fair comparison.. For nvBench, we use the EX metric for evaluation, which measures the equivalence of the generated visualizations with the ground truth based on the presented data values and chart types [52]. For VisEval, we use the Pass Rate metric to measure the ratio of valid or legal results divided by all queries, and the Readability Score judged by GPT-4V(ision) [53] to measure the overall quality of the generated visualizations [50].
@@ VII-A2 Results
As shown in Table I, DataLab achieves comparable performance with the SOTA LLM-based baselines on all four BI tasks. These baselines primarily focus on single tasks and exhibit the issues discussed in Section I, such as lacking domain knowledge incorporation or unstructured communication. Specifically, DataLab outperforms all baselines on benchmarks including DS-1000, DSEval, InsightBench, and VisEval, spreading over each critical BI stage. While certain baselines excel in individual tasks (e.g., NL2SQL), the primary focus of DataLab is to unify the BI workflow with a single LLM-based framework, maintaining satisfactory performance across various tasks. This unification is particularly beneficial for real-world scenarios requiring multi-task coordination. Moreover, many agent-based SOTA approaches (e.g., Data Interpreter [9]) that follow the reasoning and acting (ReAct) paradigm [54] can be integrated into our framework through DAG-based agent workflows (Section III). This extensibility further enhances DataLab’s practical applicability.
For tasks that require the generation of symbolic languages (e.g., NL2DSCode, NL2VIS), DataLab consistently performs well primarily due to the intermediate DSL specifications generated by our Domain Knowledge Incorporation module. Although most research benchmarks lack the necessary information for extracting table/column knowledge, DataLab adopts a meticulously-designed data profiling strategy as an alternative (Section IV-C) to fully utilize the provided data schema, enabling LLMs to correctly identify and associate the semantic relationships between data columns and NL queries, which are crucial to generate high-quality DSLs. Consequently, compared to merely feeding the original pure NL queries, these intermediate DSLs can significantly improve LLM-based agents’ performance on generating higher-level languages like SQL queries, Python code, or Vega-Lite specifications, as also shown in previous works [55].
@@ VII-D Effect of Inter-Agent Communication
TABLE II: Ablation study on Domain Knowledge Incorporation
Task / Metric S1 S2 S3
Schema Linking / Recall @5 (%) 41.02 71.79 79.49
NL2DSL / Accuracy (%) 32.52 61.66 91.10
@@ VII-E1 DAG Construction
VII-E2 Task Completion
As illustrated in Table IV, DataLab achieves a satisfactory Accuracy under both settings. Further analysis reveals that certain Markdown cells may contain critical information for task completion, which are occasionally failed to retrieve by our context retrieval mechanism due to limitations of embedding similarity [61]. This accounts for the slight 4.67% drop in Accuracy under S2. However, S2 significantly reduces the Token Cost per Query by 61.65% compared to S1, saving approximately \$65.90 for every 1,000 queries using GPT-4. This is achieved by identifying the minimum set of relevant cells based on DAGs. These results demonstrate this module’s cost-effectiveness while maintaining acceptable performance.
VIII Real-World Applications
DataLab has been effectively deployed on Tencent TEG’s Big Data Platform, achieving an average of 2,093 monthly registered users, 10,900 monthly API calls, and 319 weekly active users over a three-month period. It significantly improves the efficiency of data professionals by integrating LLM-powered BI tasks into a unified notebook interface. Below, we highlight three practical use cases to illustrate its benefits.
  • Fiona is a data engineer who regularly prepares data for product managers. Previously, she typically adds domain knowledge to prompts manually to enhance performance. With DataLab’s Domain Knowledge Incorporation module, she is surprised to discover that it can inherently understand her requirements even with ambiguous column names.
  • Henry, a data scientist specializing in recommendation algorithms, values DataLab for generating and displaying Python code directly in notebook cells. It makes editing easier and eliminates the need to repetitively copy-paste between chat-based LLM interfaces and traditional notebooks.
  • Jerry is a data analyst who provides data reports for stakeholders. Using DataLab, he cuts the time to create data visualizations by 87% (from 15 minutes to 2 minutes). He highly appreciates DataLab’s Chart cells, which enable GUI-based customization of automatically generated visualizations.
Additionally, users agree that working on DataLab facilitates communication efficiency with colleagues. For example, they can leave comments next to cells to quickly notify collaborators about their progress, avoiding the need to tediously share screenshots or code snippets as they did previously.
IX Related Work
Business Intelligence Platforms. BI platforms support users in analyzing business data for decision-making [8, 12]. Representatives like Tableau [62], Power BI [63], and Databricks [64] provide GUIs to support user interactions for data transformation and dashboard generation. These platforms also integrate natural language interfaces [65, 66] to lower the burden of manual operation. Quamar et al. [1] proposed an ontology-based method based on business models to provide semantic information and reasoning capability for query interpretation. The emergence of LLMs has further enhanced the domain knowledge integration and visualization generation abilities of BI platforms [67, 68]. Unlike existing tools that focus on visualization (e.g., Tableau, Power BI) or are optimized for engineering workflows (e.g., Databricks), DataLab provides a unified platform to satisfy various BI stages (i.e., data preparation, analysis, and visualization) and data roles (i.e., data engineers, scientists, and analysts) through a one-stop LLM-based agent framework.